OpenStudy (anonymous):

5.4 = a - (-1.8) a =

OpenStudy (xapproachesinfinity):

did you do anything?

OpenStudy (xapproachesinfinity):

what did you try

OpenStudy (anonymous):

i tryied adding it

OpenStudy (xapproachesinfinity):

can you show me what you did?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[5.4+-1.8\]

OpenStudy (anonymous):

im not good at math soooooo ya but i tryied

OpenStudy (xapproachesinfinity):

eh that's not how you do it! so first start by taking care of the two minus signs in the right hand side of the equation

OpenStudy (xapproachesinfinity):

you have to know that - times - give us +

OpenStudy (xapproachesinfinity):

so we should have \[5.4=a+1.8\]

OpenStudy (xapproachesinfinity):

got that!?

OpenStudy (anonymous):

yup

OpenStudy (xapproachesinfinity):

next is isolating a by itself on one side of the equation and the way to do that is to subtract 1.8 from both sides so \[5.4\color{red}{-1.8}=a+1.8\color{red}{-1.8}\]

OpenStudy (xapproachesinfinity):

1.8-1.8 is zero we are left with a by it self and that's what we wanted \[5.4-1.8=a\]

OpenStudy (xapproachesinfinity):

now do the subtraction

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

don't tell me you can't do subtraction!! 5.4-1.8=?

OpenStudy (anonymous):

3.6

OpenStudy (xapproachesinfinity):

correct!

OpenStudy (xapproachesinfinity):

so a=3.8

OpenStudy (anonymous):

i know how to subtract

OpenStudy (xapproachesinfinity):

you see how easy that is :)

OpenStudy (xapproachesinfinity):

yes just kidding:)

OpenStudy (xapproachesinfinity):

just study hard algebra is not that hard at all :)

OpenStudy (xapproachesinfinity):

follow you logic sense

OpenStudy (anonymous):

ok thank for all your help and teaching me how to do it

OpenStudy (xapproachesinfinity):

no problem and welcome to open study new user :)

OpenStudy (anonymous):

Welcome to OPEN STUDY @animelover47 :) Just in case you want to ask another question in this subject, make sure to close this question!